# Active-Absorbing Phase Transitions in the Parallel Minority Game

### v1

@@ VIII Conclusions and Future Directions
We have presented a comprehensive numerical study of phase transitions in the Parallel Minority Game under two distinct decision paradigms. Our key findings are that instantaneous decision rules yield critical exponents ($\beta\approx 0.95$, $\delta\approx 0.5$, $\nu_{\parallel}\approx 2.0$) consistent with mean-field directed percolation universality. Threshold-based rules produce a distinct universality class characterized by $\beta\approx 0.75$ and systematic failure of DP dynamical scaling, with evidence of multi-timescale relaxation and different exponent values in supercritical ($\delta=1.0$, $\nu_{\parallel}=0.75$) and subcritical ($\delta=0.2$, $\nu_{\parallel}=1.0$) regimes. The deviation stems from the fundamental nature of thresholding, which acts as a relevant perturbation in the renormalization group sense. The robustness of this effect across observable types and threshold variants confirms the generic nature of threshold-induced universality breaking.
@@
We should also note the possible link of the model dynamics to a frustrated spin-glass like system Sherrington and Kirkpatrick (1975). In such systems, due to both ferromagnetic and anti-ferromagnetic coupling between (Ising) spins, the ground state becomes highly degenerate and difficult to find, the energy landscape being highly corrugated. The system dynamics can often lead to a local minimum rather than the global one. Similarly, in the present system, decrease in population in one site necessarily increases the same in other sites, making the ’ground state’ or optimized distribution of crowds (presumably equal population in each site) a difficult goal. The effect of such ’frustration’ in the critical exponent values studied above, and the effectiveness of heuristic techniques (see e.g., Das, Biswas, and Chakrabarti (2025)) to find the ground states here could be an important direction.
In summary, our work establishes that minimal cognitive features at the agent level can fundamentally reshape collective critical behavior, with broad implications for modeling complex adaptive systems across disciplines.

### v2

@@ V Discussions and conclusions
We have presented a comprehensive numerical study of phase transitions in the Parallel Minority Game under two distinct decision paradigms. Our key findings are that instantaneous decision rules (strategy A) yield an active-absorbing transition in terms of the fraction of agents not in the minority or the fraction of sites where over-crowding happens, with critical exponents ($\beta\approx 0.95$, $\delta\approx 0.5$, $\nu_{\parallel}\approx 2.0$) (see Fig. 1 and  2) and consistent with mean field values for such exponents. But the threshold-based rules (strategy B) produce a distinct universality class characterized by $\beta\approx 0.75$ (see Fig. 1) and systematic failure of dynamical scaling, with evidence of multi-timescale relaxation and different exponent values in supercritical ($\delta=1.0$, $\nu_{\parallel}=0.75$) (see Fig. 3) (a-b) and subcritical ($\delta=0.2$, $\nu_{\parallel}=1.0$) (see Fig. 3) (c-d) regimes. The deviation stems from the fundamental nature of thresholding, which acts as a relevant perturbation in the renormalization group sense.
@@
We should also note the possible link of the model dynamics to a frustrated spin-glass like system Sherrington and Kirkpatrick (1975). In such systems, due to both ferromagnetic and anti-ferromagnetic coupling between (Ising) spins, the ground state becomes highly degenerate and difficult to find, the energy landscape being highly corrugated. The system dynamics can often lead to a local minimum rather than the global one. Similarly, in the present system, decrease in population in one site necessarily increases the same in other sites, making the ’ground state’ or optimized distribution of crowds (presumably equal population in each site) a difficult goal. The effect of such ’frustration’ in the critical exponent values studied above, and the effectiveness of heuristic techniques (see e.g., Das et al. (2025)) to find the ground states here could be an important direction.
In conclusion, our results demonstrate that minimal cognitive features such as decision thresholds can fundamentally alter universality classes in adaptive multi-agent systems. This has several important implications.
Real-world decision makers often exhibit inertia and threshold behavior-consumers wait for significant price changes before switching brands, investors require substantial evidence before changing portfolios, and firms delay responses until profit differentials exceed certain thresholds. Our findings suggest that such micro-level cognitive features can dramatically affect macroscopic collective behavior and phase transition characteristics.
Code Availability
The simulation code for this work is available at https://github.com/AryanTyagi-Complexity/Parallel-Minority-Game